OpenStudy (anonymous):

In a triangle, the second angle is twice the measure of the first, and the third angle is three times the measure of the second. Find the measure of the third angle.

OpenStudy (anonymous):

options.

OpenStudy (anonymous):

i got 120.

OpenStudy (johnweldon1993):

120 is correct the third angle = 120 second angle = 120/3 = 40 first angle = 40 / 2 = 20 120 + 40 + 20 = 180? Yes so 120 is correct

OpenStudy (anonymous):

ty!! :D

OpenStudy (johnweldon1993):

No problem! Good work!!